I'm trying to solve this simultaneous equation, but I can't get "x" on its own. Can you help me please?
3x + 2y = 17
2x - y = 2

Answer by edjones(8007)

3x + 2y = 17
2x - y = 2
multiply the bottom equation times 2. 4x-2y=4
now add the 2 equations vertically
3x+2y=17
4x-2y=4
---------
7x=21
divide both sides by 7.
x=3
replace x with 3 in either of the original equation.
(2*3)-y=2
6-y=2
subtract 6 from both sides. -y=-4
y=4
Check in the top equation. (3*3)+(2*4)=17
Ed Jones